# Rapid Detection of High-impedance Arc Faults in Medium Voltage Electrical Distribution Systems

### v1

@@ 4 Results and Discussion
The following section provides the findings from the cases conducted to evaluate the effectiveness of the arc fault detection framework. The analysis employs the system illustrated in 1. The simulation integrates an arc model derived from [19], which serves as a fundamental reference for the construction and use of the model in the study. In our study, arc faults were strategically placed at different locations between Bus 1 and Bus 2, as well as between Bus 2 and Bus 3, within the medium voltage distribution system. This placement was done to examine the detection performance across different regions of the network. Altering the fault locations offered an evaluation of the robustness and precision of the arc fault detection system in recognising faults at different points along the distribution line. This approach ensures a full evaluation of the detection system over a range of possible fault conditions.
All the studies have included the data containing both arc fault and normal current waveforms. As demonstrated in the arc model, variations in the model’s variable aspects, particularly DURATION, EXTENT, and OFFSET, result in unique waveforms. The values were set based on an existing reference for a 50 Hz system with a phase-voltage level of 12 kV (the maximum voltage) for our system. The parameter ranges are as follows: $\text{OFFSET}\in(-12,12)\,\text{kV},\,\text{EXTENT}\in(0,+\infty)\,\text{k}\Omega,\,\text{DURATION}\in(0,20)\,\text{ms}$ , which align with typical fault current behavior in medium voltage systems [19]. The selected values guarantee that the arc fault current graph aligns accurately with the graphs presented in [19], with the specific values utilised in each case detailed in Table 1 for reference.
@@ 4.3 Case C: Arcing Fault Present when grounded in wet cement
In practical HIAFs, fault current distortions are greatly affected by multiple parameters, especially surface humidity conditions [34]. This study investigates HIAF under conditions when it is grounded on wet cement, with DURATION set to 0.007 seconds and Extent set to 50000 $\Omega$ OFFSET set to 0.2 kV. A simulation was performed for 0.5 seconds, with the arc fault initiated at 0.2 seconds. The subsequent deviation of the forcing operator was recognized at 0.2001 seconds, proving the system’s capabilities to identify the arc fault within a short interval of 0.1 milliseconds. The value of the forcing operator in this case was calculated to be 0.15201, which lies inside the threshold range for arc faults. This scenario is represented in Figure 9.
@@ 4.5 Case E: Nonarcing disturbance (Load Switching)
Figure 10: Non arcing disturbance (load switching).
@@ 4.7 Case G: Arcing fault current with Induction Motor Load
In this scenario, Load 1 stays constant at 10 MW, while Load 2 comprises an induction motor (IM) operating at 3.3 kV with a rated capacity of 0.2 MVA. The motor is driven through a step-down transformer, which reduces the voltage from 12 kV to 3.3 kV. The simulation runs for 0.5 seconds, with the fault introduced at 0.2 seconds and continuing for 0.1 seconds. Two scenarios have been investigated involving an IM load: one with a low-current arcing fault and the other with a high-current arcing fault. These scenarios are represented in Figures 13 and 13, respectively. In the low-current arcing fault scenario, a deviation in the forcing operator was noted at 0.20015 seconds, with the value of the forcing operator measured to be -0.08045. In the high-current arcing fault scenario, the deviation visible at 0.2001 seconds, with the value of forcing operator -0.08046. Both values fall within the threshold range for arc faults, confirming the robustness and effectiveness of our approach.
@@ 4.8 Case H: Impact of Noise
To evaluate the influence of measurement noise, white Gaussian noise is introduced to the current measurements. Simulations were executed with signal-to-noise ratios (SNR) of 60 dB, 70 dB, and above, during which our approach exhibited excellent performance. Incorporating white Gaussian noise at this level and higher guarantees a realistic simulation of noise that distorts the signal without substantially compromising its integrity, thus allowing an appropriate evaluation of the method’s effectiveness under realistic circumstances. In cases of extremely low signal-to-noise ratio, denoising filters may be employed for efficient noise suppression. Figure 15 depicts a scenario with a signal-to-noise ratio of 70 dB. The deviation of the forcing operator is noted at 0.20065 seconds, with a value of 0.116793459. In presence of noise, it has been found that the detection time is extended. However, the detection is still robust from the performance of our analysis.
@@
Cases Extent ($\Omega$) Duration (seconds) Offset (kV) $R_{T}$ ($\Omega$) Forcing Operator Time of Detection (ms)
A 5000 0.00413 0.2 1000 -0.1035 0.45
B 5000 0.00413 0.2 0.001 -0.0776 0.05
C 50000 0.00700 0.2 50 0.1520 0.10
D 4708 0.00700 0.2 50 -0.1614 0.15
E NA NA NA NA 0.0429 0.20
F NA NA NA NA -0.2020 0.10
G (a) 5000 0.00413 0.2 1000 -0.0805 0.15
G (b) 5000 0.00413 0.2 0.001 -0.0805 0.10
H 5000 0.00413 0.2 1000 0.1168 0.65
Table 1 provides an overview of the analysis of the experiments, including the corresponding time of detection for each conducted experiment. The range of the forcing operator permits the differentiation between arc faults and other types of faults, as arc faults present a lower range and require more time for detection compared to other fault types. It has been shown that a forcing operator range of +0.06 to + 0.18 and -0.06 to -0.18 helps the rapid identification of arc faults. In contrast, when the forcing operator range changes, such as +0.2 and higher values and -0.2 and lower values, the system identify other faults, while non-arcing disturbances, such as load switching, are identified with a forcing operator range of +0.045 to -0.045.
@@ 5 Comparison With Existing Techniques
A comparative analysis has been conducted between the proposed method and existing arcing fault detection methods found in the literature. The standard detection time for high-impedance arc faults in medium voltage networks may vary based on the specific detection methods and technologies employed. Generally, high-impedance arc faults are hard to detect because of their low fault current and stochastic nature, which can result in longer detection times compared to other fault types. Usually, detection times for high-impedance faults vary between milliseconds to several seconds. Several approaches available in the literature demonstrate varied detection times for HIFs. The Mathematical Morphology technique [35] detects HIFs in 80 milliseconds, while the Harmonic Content-based method[36] yields a detection time of 200 milliseconds. Additionally, the SVM and DWT based [37] technique provides a substantially faster detection time of 0.90 milliseconds. Improvements in technology continue to push for faster and more reliable detection systems in medium voltage networks. However, our proposed method demonstrates the early detection capability of arc faults.A comprehensive literature study comparing the proposed methodology with existing methodologies, demonstrating that the proposed method is 50% more faster in arc fault detection. The findings are displayed in Table 2. It is evident that the proposed method exhibits the best level of accuracy in detecting arcing faults.
Table 2: Comparison With Existing Techniques.
Techniques Detection time (ms) Detection rate %
Mathematical Morphology technique [35] 80 100
Harmonic Content-based method[36] 200 -
SVM and DWT based technique [37] 0.90 97.6
Proposed method 0.45 100
@@ 6 Conclusion
Our research greatly contributes to the field of arc fault detection by addressing crucial gaps in existing knowledge. The application of data obtained by PSCAD simulations of medium voltage distribution lines distinguishes our analysis, filling a noteworthy gap in earlier research largely focused on low-voltage systems. The implementation of HAVOK analysis, derived from nonlinear dynamics, constitutes a shift from state-of-the-art black-box feature-based approaches. The selection not only offers interpretable features but also exposes innovative application possibilities in the context of arc fault detection.
The main achievement of the suggested technique is its capacity to identify arc faults in a minimal duration, with detection times as minimal as 0.45 ms and 0.05 ms. This rapid detection ability markedly improves the system’s dependability and reaction time, mitigating the chance of damage and enhancing the overall safety of the distribution network. Future work focuses on confirming the simulation results with real-world data. Conducting field tests would provide a more comprehensive understanding of the system’s performance and potential obstacles in real applications.
The insights provided here pave the path for future developments in the domain, promoting a more secure and trustworthy operational environment.

### v2

@@ 4 Results and Discussion
The following section provides the findings from the cases conducted to evaluate the effectiveness of the arc fault detection framework. The analysis employs the system illustrated in 1. The simulation integrates an arc model derived from [24], which serves as a fundamental reference for the construction and use of the model in the study. In our study, arc faults were strategically placed at different locations between Bus 1 and Bus 2, as well as between Bus 2 and Bus 3, within the medium voltage distribution system. This placement was done to examine the detection performance across different regions of the network. Altering the fault locations offered an evaluation of the robustness and precision of the arc fault detection system in recognising faults at different points along the distribution line. This approach ensures a full evaluation of the detection system over a range of possible fault conditions. Simulations were carried out at a sample rate of 20 kHz (i.e., 50 µs resolution) to ensure exact time-domain analysis. The algorithm was executed in Python on a normal workstation equipped with an Intel i5 CPU at 2.60 GHz and 16 GB of RAM.
All the studies have included data containing both arc fault and normal current waveforms. As demonstrated in the arc model, variations in the model’s variable aspects, particularly DURATION, EXTENT, and OFFSET, result in unique waveforms. The values were set based on an existing reference for a 50 Hz system with a phase-voltage level of 12 kV (the maximum voltage) for our system. The parameter ranges are as follows: $\text{OFFSET}\in(-12,12)\,\text{kV},\,\text{EXTENT}\in(0,+\infty)\,\text{k}\Omega,\,\text{DURATION}\in(0,20)\,\text{ms}$ , which aligns with typical fault current behavior in medium voltage systems [24]. The selected values guarantee that the arc fault current graph aligns accurately with the graphs presented in [24], with the specific values utilised in each case detailed in Table 2 for reference.
@@ 4.3 Case C: Arcing Fault Present when grounded in wet cement
In practical HIAFs, fault current distortions are greatly affected by multiple parameters, especially surface humidity conditions [36]. This study investigates HIAF under conditions when it is grounded on wet cement, with DURATION set to 0.007 seconds and Extent set to 50000 $\Omega$, OFFSET set to 0.2 kV. A simulation was performed for 0.5 seconds, with the arc fault initiated at 0.2 seconds. The subsequent deviation of the forcing operator was recognized at 0.2001 seconds, proving the system’s capabilities to identify the arc fault within a short interval of 0.1 milliseconds. The value of the forcing operator in this case was calculated to be 0.15201, which lies inside the threshold range for arc faults. This scenario is represented in Figure 9.
@@ 4.5 Case E: Nonarcing disturbance (Load Switching)
Figure 10: Non-arcing disturbance (load switching).
@@ 4.7 Case G: Arcing fault current with Induction Motor Load
In this scenario, Load 1 stays constant at 10 MW, while Load 2 comprises an induction motor (IM) operating at 3.3 kV with a rated capacity of 0.2 MVA. The motor is driven through a step-down transformer, which reduces the voltage from 12 kV to 3.3 kV. The simulation runs for 0.5 seconds, with the fault introduced at 0.2 seconds and continuing for 0.1 seconds. Two scenarios have been investigated involving an IM load: one with a low-current arcing fault and the other with a high-current arcing fault. These scenarios are represented in Figures 13 and 13, respectively. In the low-current arcing fault scenario, a deviation in the forcing operator was noted at 0.20015 seconds, with the value of the forcing operator measured to be -0.08045. In the high-current arcing fault scenario, the deviation is visible at 0.2001 seconds, with the value of the forcing operator -0.08046. Both values fall within the threshold range for arc faults, confirming the robustness and effectiveness of our approach.
@@ 4.8 Case H: Impact of Noise
To evaluate the influence of measurement noise, white Gaussian noise is introduced to the current measurements. Simulations were executed with signal-to-noise ratios (SNR) of 60 dB, 70 dB, and above, during which our approach exhibited excellent performance. Incorporating white Gaussian noise at this level and higher guarantees a realistic simulation of noise that distorts the signal without substantially compromising its integrity, thus allowing an appropriate evaluation of the method’s effectiveness under realistic circumstances. In cases of extremely low signal-to-noise ratio, denoising filters may be employed for efficient noise suppression. Figure 15 depicts a scenario with a signal-to-noise ratio of 70 dB. The deviation of the forcing operator is noted at 0.20065 seconds, with a value of 0.116793459. In the presence of noise, it has been found that the detection time is extended. However, the detection is still robust from the performance of our analysis.
@@
Cases Extent ($\Omega$) Duration (seconds) Offset (kV) $R_{T}$ ($\Omega$) Forcing Operator Time of Detection (ms) Arc/Disturbance Type
A 5000 0.00413 0.2 1000 -0.1035 0.45 LCAF
B 5000 0.00413 0.2 0.001 -0.0776 0.05 HCAF
C 50000 0.00700 0.2 50 0.1520 0.10 ACWC
D 4708 0.00700 0.2 50 -0.1614 0.15 ACDS
E NA NA NA NA 0.0429 0.20 NAD (LS)
F NA NA NA NA -0.2020 0.10 LGF
G (a) 5000 0.00413 0.2 1000 -0.0805 0.15 LCAFIM
G (b) 5000 0.00413 0.2 0.001 -0.0805 0.10 HCAFIM
H 5000 0.00413 0.2 1000 0.1168 0.65 AFWN
Notation: LCAF — Low Current Arc Fault, HCAF — High Current Arc Fault, NAD (LS) — Non-Arcing Disturbance (Load Switching), LGF — Line to ground fault , LCAFIM — Low Current Arc Fault with Induction motor load, HCAFIM — High Current Arc Fault with Induction motor load, AFWN — Arc Fault with Noise.
Table 2 provides an overview of the analysis of the experiments, including the corresponding time of detection for each conducted experiment. The range of the forcing operator permits the differentiation between arc faults and other types of faults, as arc faults present a lower range and require more time for detection compared to other fault types. It has been shown that a forcing operator range of +0.06 to + 0.18 and -0.06 to -0.18 helps the rapid identification of arc faults. In contrast, when the forcing operator range changes, such as +0.2 and higher values and -0.2 and lower values, the system identify other faults, while non-arcing disturbances, such as load switching, are identified with a forcing operator range of +0.045 to -0.045.To assess detection robustness, we conducted 50 simulations for each scenario. The proposed technique attained a detection accuracy of 100% for arc faults, with 0% false positives for non-arcing disturbances. The detection exhibited consistency across the trials, proving its robustness.
@@ 5 Comparison With Existing Techniques
A comparative analysis has been conducted between the proposed method and existing arcing fault detection method found in the literature. The standard detection time for high-impedance arc faults in medium voltage networks may vary based on the specific detection methods and technologies employed. Generally, high-impedance arc faults are hard to detect because of their low fault current and stochastic nature, which can result in longer detection times compared to other fault types. Usually, detection times for high-impedance faults vary between milliseconds to several seconds.Our proposed method is carefully assessed against the method outlined in [37], employing the same dataset for a direct and thorough comparative study. This method in [37] identifies arc faults by examining waveform features through two approaches: it employs the Fast Fourier Transform (FFT) to measure harmonic unpredictability, signifying unstable arcing, and the Discrete Wavelet Transform (DWT) to evaluate waveform distortion, indicating stable arcing. The properties are assessed against thresholds established from a healthy signal baseline to ascertain the occurrence and duration of arc faults.The technique detects HIAFs in 70 milliseconds. Improvements in technology continue to push for faster and more reliable detection systems in medium voltage networks. However, our proposed method demonstrates the early detection capability of arc faults. The findings, as outlined in Table 3, indicate that the suggested method attains a detection time of 0.45 ms, which is 99.36% enhanced than the Harmonic Randomness and Waveform Distortion approach (70 ms) and exceeds the performance of other  techniques by greater margins.
The findings demonstrate that the suggested method is both efficient and resilient under diverse demanding settings, including low and high current arc faults, different grounding materials, and noisy environments. The method’s dependence on Koopman-invariant forcing coordinates enables it to accurately differentiate arc faults from non-arcing disturbances, a task that many black-box AI methodologies find challenging. The method’s reliable performance, even amongst measurement noise (with an SNR as low as 60 dB), underscores its practical feasibility.
Moreover, the physics-informed foundation of the HAVOK model offers interpretable features, which is a significant benefit in safety-critical fields where explainability is as vital as precision. This study establishes a solid basis for future arc fault detection systems in intricate electrical networks by the integration of rapid detection, resilience, and interpretability.
Table 3: Comparative Summary of Existing Arc Fault Detection Techniques and Advancements by the Proposed Method
Method Technique Detection Time Limitations Advancement by This Work
Harmonic Randomness and Waveform Distortion[37] Harmonic analysis of current signals 70 ms Conventional harmonic and DWT techniques may exhibit sensitivity to fluctuating load circumstances, noise, and non-arc disturbances, which could result in false positives or undetected anomalies without careful calibration and a solid baseline foundation. 99.36% enhaced detection
Proposed Method Koopman operator-based HAVOK with forcing signature 0.45 ms Requires parameter tuning for unseen networks Fast, explainable detection using Koopman-invariant forcing operator; resilient under noise and non-ideal conditions
@@ 6 Conclusion
Our research greatly contributes to the field of arc fault detection by addressing crucial gaps in existing knowledge. The application of data obtained by PSCAD simulations of medium voltage distribution lines distinguishes our analysis, filling a noteworthy gap in earlier research largely focused on low-voltage systems. This study proposed a unique, physics-informed methodology for detecting high-impedance arc faults (HIAFs) in MV electrical distribution systems using the HAVOK analysis. By integrating Koopman operator theory into nonlinear current dynamics, the method made major improvements in detection speed and interpretability over traditional and black-box AI-based techniques. The suggested approach successfully detects arc faults under a wide range of situations, including different grounding surfaces, noise levels, and load types, achieving detection within as little as 0.45 milliseconds. Comparative investigation indicated that the strategy is 99.36% enhanced over other existing methods, while maintaining good detection accuracy. These contributions boost the reliability and responsiveness of arc fault detection systems, representing a considerable advancement in the field of power system safety.
We acknowledge that the present study is confined to simulation-based validation. Field testing of medium-voltage systems presents practical and safety concerns; future efforts will incorporate hardware-in-the-loop (HIL) configurations and experimental validation with lab-scale prototypes to further substantiate the suggested detection method. For future work, the proposed approach should be tested using real-world datasets obtained from physical testbeds or operational distribution networks to ensure its robustness outside of simulation. The insights provided here pave the path for future developments in the domain, promoting a more secure and trustworthy operational environment.